I am in need of a geometry refresher, so, whoever is kind enough to answer this, please explain as to how you derived your answer. What is the length of each diagonal of a rectangular box with length 53 cm, width 48 cm, and height 70 cm?
Answer by scott8148(6628)

the length of the diagonal is sqrt[(length)^2+(width)^2+(height)^2]

this is a serial application of the Pythagorean theorem

if you take the diagonal of any face, it is the hypotenuse of a triangle that includes two of the dimensions
__ eg. d^2=L^2+W^2

the diagonal of the box, is the hypotenuse of a triangle that includes the "face" diagonal and the 3rd dimension
__ eg. D^2=d^2+H^2 __ D^2=L^2+W^2+H^2